OpenStudy (anonymous):

Let f(x) = 4x + 3 and g(x) = -2x + 5. Find (f)(g)(5) a. 23 b. –41 c. –17 d. –5

OpenStudy (anonymous):

make x=5 in both f(x) and g(x) then multiply them that's it do it

OpenStudy (anonymous):

thats where im confused at, the actual multipling both of them , can u show me ..?

OpenStudy (jdoe0001):

multiply, the same way you would if you were dealing with two digits pick only 1 term, multiple times all others on the other group pick another term, and do the same then simplify

OpenStudy (anonymous):

f(x) and g(x) into (f)(g)(5) is multiplying right? You don't mean f(g(x))? If not, you multiply the equations while inputting 5 for x. 4(5) + 3 times -2(5) + 5.

OpenStudy (jdoe0001):

(a + b) * (c + d) a * c a* d then b * c b * d add them up, simplify

OpenStudy (anonymous):

nah i dont mean that , u got it , but okayy i see what yall are sayin , appreciate it

OpenStudy (jdoe0001):

\(\bf f(x) = 4x + 3 \qquad g(x) = -2x + 5 \\ \quad \\ (f)(g)(5)\implies f({\color{brown}{ 5}})\cdot g({\color{brown}{ 5}})\implies (4({\color{brown}{ 5}})+3)\cdot (-2({\color{brown}{ 5}})+5) \) even simpler